### v1

# Who Made This? Fake Detection and Source Attribution with Diffusion Features

Abstract
The rapid progress of generative diffusion models has enabled the creation of synthetic images that are increasingly difficult to distinguish from real ones, raising concerns about authenticity, copyright, and misinformation. Existing supervised detectors often struggle to generalize across unseen generators, requiring extensive labeled data and frequent retraining. We introduce FRIDA (Fake-image Recognition and source Identification via Diffusion-features Analysis), a lightweight framework that leverages internal activations from a pre-trained diffusion model for deepfake detection and source generator attribution. A k-nearest-neighbor classifier applied to diffusion features achieves state-of-the-art cross-generator performance without fine-tuning, while a compact neural model enables accurate source attribution. These results show that diffusion representations inherently encode generator-specific patterns, providing a simple and interpretable foundation for synthetic image forensics.
Index Terms —
Overview of our framework for deepfake classification and attribution. Given an input image, our Image Prototyping module extracts image prototype from the U-Net of a Stable Diffusion Model. This compact representation is then used for two downstream tasks: (i) Deepfake classification via a k-NN classifier to determine whether the image is real or fake, and (ii) Deepfake attribution via an MLP to identify the generative model (e.g., GLIDE, Stable Diffusion, BigGAN) responsible for the fake.
1 Introduction
The outstanding capabilities of recent generative models to synthesize media with exceptional fidelity and variety [1, 16, 4, 36] have accelerated the proliferation of AI‑generated content across domains, including medical imaging [23, 47, 42], education [53, 51], marketing [20, 12], and robotics [50]. This rapid progress has been largely driven by diffusion-based generators — including Stable Diffusion Model (SDM) [36], DALL-E  [35], and Imagen [37] — which produce highly realistic and semantically coherent visuals from simple text prompts. This increasing accessibility of advanced image generation tools through user-friendly interfaces poses considerable risks, including the creation of biased content and the violation of intellectual property rights. To mitigate these risks, robust detection and attribution methods are required to reliably identify synthetic media and determine their source models [39, 13, 29].
Early fake detection methods relied on supervised learning pipelines [52, 59, 29], which depend on vast labelled datasets and intensive computation. Approaches for source generator attribution pushed these requirements even further, as they must learn generator-specific fingerprints [54, 8, 46]. However, as generative models evolve rapidly, the continual need for data collection and retraining makes classifier-based detectors hard to scale, underscoring the need for lightweight, data-efficient methods that generalize across both known and emerging generators.
An opportunity to move beyond task-specific classifiers has emerged with the rise of large-scale foundation models based on Vision Transformers (ViTs) [19], e.g., CLIP [34] and DINO [9]. Pre-trained on massive datasets, these models learn general, transferable representations that have shown remarkable effectiveness in downstream tasks [24, 45] — including deepfake detection [31, 40, 44] — when their extracted features are used for classification.
Similarly, diffusion-based generative models, though designed for image generation, have proven to be remarkably effective feature encoders. Recent studies have shown that pre-trained SDMs provide rich semantic information about image content [5, 6], enabling strong performance in applications such as semantic segmentation [3, 15], and other vision-related problems — substantially reducing the computational cost of training large models from scratch. At the same time, diffusion models have been increasingly used for fake detection. Some methods identify diffusion-generated images through reconstruction error [49], while others analyse latent trajectories via diffusion inversion to enhance generalisation across unseen generators [10, 43]. Based on the principle that images from different sources exhibit distinct distributions, Zhong et al. [57] employ a SDM as a denoising tool to identify fake images by detecting the resulting artifacts in their reconstructed versions. However, these approaches typically require repeated diffusion or inversion steps, making them computationally expensive and less practical for large-scale or real-time detection scenarios.
Inspired by these recent advances, we explore whether the internal features of a pre-trained diffusion-based generative model can effectively separate real from synthetic images and, beyond detection, enable attribution to the source generator. To this end, we introduce a lightweight and data-efficient framework that operates entirely within the latent feature space of SDMs at the last inference step. Our approach detects deepfakes by applying a training-free $k$-Nearest Neighbours ($k$-NN) classifier to latent representations extracted from the SDM. This approach achieves state-of-the-art performance on the GenImage benchmark  [59] for fake image detection. This strategy eliminates the need for costly inversion or fine-tuning and shows strong generalisation to unseen generators. We further extend our method to source-model attribution, showing that a lightweight Multi-Layer Perceptron (MLP) trained on the SDM latent features can accurately identify the generator responsible for a synthetic image.
The main contributions of our work can be summarized as follows:
  1. 1.
    We show that latent features extracted from specific layers of a pre-trained SDM are highly discriminative for both fake image detection and source model attribution.
  2. 2.
    We benchmark our fake image detection approach on data generated by eight different models, achieving state-of-the-art performance with a novel, lightweight $k$-NN-based framework that is entirely training-free. Our method requires only a small support set, generalises effectively to unseen generators, and adapts to new data without any fine-tuning.
  3. 3.
    A lightweight neural classifier trained on the SDM latent representation proves to be highly effective in source model attribution, suggesting the presence of generator-specific characteristics in the features. We use SHAP (SHapley Additive exPlanations) [28] to further investigate the capability of the SDM to encode these model-specific signatures.
The remainder of this paper is organized as follows:  Section 2 reviews relevant literature on synthetic image detection and attribution.  Section 3 describes the GenImage dataset and the procedure used to extract the latent representation of each image using the SDM.  Section 4 describes the experimental setup, while  Section 5 presents and discusses the obtained results. Finally,  Section 6 concludes the paper with a summary of our findings and outlines future research directions.
2 Related Work
Frequency- and Texture-based Detectors.
Early research on detecting GAN-generated content showed that synthetic images contain characteristic frequency patterns and texture artifacts. Zhang et al. [56] introduced AutoGAN, a frequency-domain simulator of upsampling artifacts, and trained a spectrum-based classifier to detect periodic Fourier patterns without access to the generator. Similarly, Wang et al. [46] showed that CNN-generated images share common low-level artifacts across architectures and datasets. Their ResNet-based detector, trained on ProGAN, generalized remarkably well to unseen generators such as StyleGAN and BigGAN, revealing universal CNN fingerprints that persist across models. Gram-Net [26] models global texture correlations via multi-layer Gram matrices, improving robustness to compression and noise and enabling better cross-GAN generalization by leveraging long-range texture cues. Qian et al. [33] advanced this line of work with F3-Net, a two-branch frequency-aware CNN classifier combining frequency decomposition and local statistics via cross-attention, outperforming spatial-domain detectors on a deepfake faces dataset even under strong compression.
Diffusion-based Detection and Reconstruction Methods.
Following the success of diffusion-based models, several works have focused on detecting diffusion-generated images. DE-FAKE [38] is a hybrid two-branch architecture that exploits image-text consistency between captions and visual content to detect synthetic images. DIRE [49] detects diffusion-generated images by inverting an input into the diffusion latent space and then reconstructing it through the full reverse denoising trajectory; the discrepancy between the original and reconstructed signals is used for detection. Similarly, LaRE2 [29] estimates a Latent Reconstruction Error using only a single-step reconstruction in latent space, using a module that aligns and refines features across spatial and channel dimensions. A related line of work [48] also leverages a pre-trained diffusion model, extracting multi-timestep responses under the hypothesis that synthetic images, which lie outside the natural image manifold, exhibit distinctive denoising behavior. ESIDE [52] incorporates frequency perturbations into diffusion inversion, training an ensemble of CLIP-based classifiers on noised representations to improve robustness and interpretability. More recently, LATTE [43] introduces a Latent Trajectory Embedding framework that explicitly models the temporal evolution of latent representations across denoising steps. By aggregating multi-timestep latent features through joint visual–latent refinement, LATTE captures dynamic generation cues beyond reconstruction error, achieving strong cross-generator and cross-dataset generalization. An alternative approach [57] treats detection as an anomaly detection task. They learn the low-level feature distribution of real images by training an extractor to spot pixel-level differences between original images and their denoised counterparts, effectively identifying generated content that falls outside this learned distribution.
Semantic and Open-Set Attribution Frameworks.
Recent research explores semantic alignment and open-set detection to generalize across diverse generators. Zhu et al. [58] propose MAID, a framework-agnostic attribution method that extracts Diffusion Model Activations (DMA) by treating pre-trained diffusion models as denoising autoencoders. These activations encode model-specific patterns without requiring white-box access or prompts, supporting both detection and attribution. SemGIR [55] employs semantic-guided image regeneration: a candidate image is captioned, regenerated via text-to-image synthesis, and compared with its reconstruction using CLIP-based encoders. This forces the detector to focus on generator-specific artifacts rather than prompt semantics and achieves strong cross-generator generalization. Cioni et al. [14] move beyond frequency-based fingerprints by leveraging intermediate representations of large ViT-based models such as CLIP and DINOv2. Their open-set attribution framework combines linear probing and k-nearest neighbors with confidence- or distance-based rejection to identify images from unseen generators, providing a unified and retraining-free solution across GAN and diffusion sources.
3 Materials and Methods
3.1 GenImage Dataset
GenImage [59] is a large-scale benchmark designed to detect generated imagery, composed of both synthetic and real images sourced from ImageNet [17]. The dataset contains roughly 1.35 million synthetic images produced by eight distinct generative models: BigGAN [7], GLIDE [32], VQDM [21], SDM (v1.4 and v1.5)[36], ADM[18], Midjourney [1], and Wukong [2]. Both the real and synthetic images cover the 1,000 ImageNet classes, with the synthetic portion providing approximately 1,350 images per class (1,300 for training and 50 for testing). The generated images were created using simple text prompts following the template “photo of [class]” and have resolutions ranging from $128\times 128$ to $1024\times 1024$, depending on the source model. In this study, we utilize a subset of the training images. For each generator, we randomly sampled 10,000 real and 10,000 synthetic images (10 images per ImageNet class). These subsets are then partitioned into training ($80\%$) and validation ($20\%$) sets. In the rest of the paper, we will refer to these two sets as “training subset” and “validation set”. Finally, the original GenImage test set was used as a held-out evaluation set.
3.2 Image Prototype Extraction
We exploit a pre-trained SDM v1.5111https://github.com/hkproj/pytorch-stable-diffusion and we use the following procedure for extracting internal features prototype from a specific layer (see Figure 1).
Figure 1: Prototype extraction from Stable Diffusion U-Net. In this example, we extract and average the features from the first decoder layer at $16\times 16$ resolution.
Each input image is resized to $512\times 512$, and we run a forward pass at $t=0$, which corresponds to the final denoising step of the diffusion U-Net. After encoding the image into a latent representation via the VAE, this latent is passed to the U-Net. Compact prototypes are then obtained by spatially averaging the feature maps extracted from a chosen U-Net layer.
4 Experimental Setup
This section details the experimental setup used in our study. First, we compare features from different layers of the diffusion U-Net to identify which layer is most effective at distinguishing real from synthetic content. We then describe the experimental setups for our two primary tasks: fake image detection (Section 4.2) and source generator attribution (Section 4.3).
4.1 U-Net Layer Selection by Linear Probing
To identify which layers provide the most effective features for deepfake detection, we extract latent representations from the encoder, bottleneck, and decoder layers of the U-Net at multiple spatial resolutions ($64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$). Following the GenImage protocol [59], we train eight classifiers, one for each generator in the dataset, and evaluate each classifier on images from all other generators. In particular, we adopt a linear probing setup — commonly used to evaluate the quality and separability of learned representations in self-supervised and foundation models  [11, 22, 34] — by training a linear classifier with sigmoid activation on the latent prototypes from each layer. Each classifier is trained on the training subset (see Section 3.1) with real and synthetic images from a given generator. All models are trained with the AdamW optimizer [27] using a learning rate of $3\times 10^{-4}$. Training is early stopped based on validation accuracy, with a patience of 15 epochs. To assess generalization, we adopt a cross-generator evaluation: following common practice [59], each classifier trained on a specific generator is evaluated on the fake detection task using its own validation set (seen data distribution) as well as the validation sets of the seven remaining generators (unseen data distributions). The U-Net layer whose features yield the highest average cross-generator accuracy is then selected and used for all subsequent experiments.
4.2 Fake Image Detection
Our first goal is to design a fake image detection pipeline capable of generalizing to images from unseen generators. To this aim, we evaluate two approaches: a simple neural network (MLP) and a distance-based classifier ($k$-NN). In both cases, we follow the cross-generator evaluation protocol described in Section 4.1, using the selected U-Net layer.
MLP for Fake Image Detection.
We evaluate three MLPs with different numbers of hidden units and hidden layers. In particular, we consider the following configurations:
  • MLP-640: A single hidden layer with 640 units.
  • MLP-320: A single hidden layer with 320 units.
  • MLP-640-320: Two hidden layers with 640 and 320 units, respectively.
All the models have two output neurons with a softmax activation function and are trained using binary cross-entropy loss with AdamW optimizer and a learning rate of $3\times 10^{-4}$. The validation set accuracy is used to early stop the training process with a patience of 15 epochs.
$k$-NN for Fake Image Detection.
[…]
The best MLP and $k$-NN models are selected following the cross-generator protocol described in Section 4.1, based on validation accuracy, and then tested on the GenImage test set.
4.3 Source Model Attribution
Our second objective is source generator attribution, a multi-class classification task to identify the source model of an image. As for fake image detection, we compare the MLP and $k$-NN approaches, adapting the models and the evaluation protocol for multi-class classification. More specifically, we employ a single classifier with nine output classes (eight generators and real images).
MLP for Source Model Attribution.
We evaluate the same MLP configurations used for fake-image detection (see Section 4.2), but replace the output layer with a 9-way softmax classifier and train with cross-entropy loss. All MLPs are trained on synthetic images from the training subsets of all generators, while real images are sampled from the training subset of the Midjourney dataset. All models are trained using the AdamW optimizer with a learning rate of $3\times 10^{-4}$, and the validation accuracy is used to early stop the training process using a patience of 15 epochs.
$k$-NN for Source Model Attribution.
[…]
The most effective approach is then evaluated on the GenImage test set for the attribution task.
5 Results
This section presents our experimental results, starting with the analysis of features extracted from the diffusion U-Net (Section 5.1), and then covering fake image detection (Section 5.2) and source model attribution (Section 5.3).
5.1 U-Net Layer Selection by Linear Probing
Following the experimental setup in Section 4.1, we test image prototypes from all the layers of the diffusion U-Net using a linear probing approach to determine (i) if the latent space representation can be used to distinguish between real and synthetic data, and (ii) which layer produces the most informative representation for this task. Figure 2 plots the average cross-generator validation accuracy for features extracted from different levels of the diffusion U-Net.
Figure 2: U-Net Layer Selection by Linear Probing. Average cross-generator validation accuracy obtained using the prototypes extracted from different U-Net layers (labelled as Encoder, Bottleneck or Decoder, followed by the spatial resolution and the intra-stage index). The best accuracy is achieved by features from the first layer of the decoder at $16\times 16$ resolution.
The results of our feature-probing analysis show that the latent features extracted from the diffusion U-Net can be highly discriminative for distinguishing real from synthetic images. In particular, we identify the first layer of the decoder at $16\times 16$ resolution (Decoder 16_0) as the source of the most informative features, yielding a peak average cross-generator validation accuracy of $78.50\%$. Given this result, we adopt these features in the remainder of our study.
5.2 Fake Image Detection
Following the experimental setup described in Section 4.2, we evaluate two methods: one based on neural networks (i.e., MLP) and the other on a $k$-NN approach.
Neural Network Classifier.
We conduct preliminary experiments using the three MLPs described in Section 4.2, MLP-320, which achieves an average cross-generator accuracy of $79.17\%$, outperforms MLP-640 ($78.23\%$) and MLP640-320 ($78.94\%$). A detailed breakdown of the results for MLP-320 on cross-generator fake image detection is presented in Table 1.
Trained on Tested on
Midj. SDV1.4 SDV1.5 ADM Glide Wukong VQDM BigGAN Avg.
Midj. 98.0 85.6 84.6 68.0 89.9 79.7 52.1 49.8 75.9
SDV1.4 89.2 98.8 98.5 61.7 76.3 94.5 61.2 49.1 78.6
SDV1.5 87.5 98.8 98.8 59.1 70.9 94.8 60.5 49.3 77.5
ADM 67.6 74.2 75.5 99.3 98.8 68.1 88.3 93.9 83.2
Glide 63.8 64.7 64.9 94.4 99.7 59.1 85.3 87.7 77.5
Wukong 88.3 98.0 98.2 86.6 93.6 98.1 86.6 59.2 88.6
VQDM 57.6 64.4 64.7 94.9 99.5 61.4 99.7 99.5 80.2
BigGAN 54.0 53.0 53.0 79.8 99.4 53.1 80.2 99.8 71.5
Avg. 75.7 79.7 79.8 80.4 91.0 76.1 76.7 73.5 79.1
Table 1: Cross-generator evaluation of the MLP-320 for fake image detection. We train the model on each generator and test it on the validation set of all the generators.
The results reveal a clear pattern: while the MLP effectively learns generator-specific artifacts (as evidenced by the high diagonal accuracy), it struggles to generalize to unseen generators. This behavior suggests the presence of distinct generator families that share underlying characteristics within the latent space of the diffusion U-Net. For instance, the MLP trained on Midjourney data performs well on images from SDV1.4, SDV1.5, and Glide, indicating a latent similarity among their generated outputs.
$k$-NN Classifier.
While the features extracted from the diffusion U-Net are discriminative enough to separate real and synthetic data, the MLP classifier tends to learn specific artifacts of a given generator, which limits its ability to generalize. To address this limitation, we propose a $k$-NN approach. As a training-free method that relies on distances between feature prototypes, we hypothesize that $k$-NN can achieve a more robust generalization than the MLP.
First, we identify the optimal $k$-NN configuration following the extensive parameter search described in Section 4.2. The model using the correlation distance, with $k=101$ and a support size of $2000$, achieved the best performance on the validation set (see Table 2). It yielded an average accuracy of $85.3\%$ across the eight generators, a significant improvement over the $79.17\%$ obtained by the best MLP approach. This optimized version of the $k$-NN model is finally tested on the GenImage test set (Table 3).
Support set from Tested on
Midj. SDV1.4 SDV1.5 ADM Glide Wukong VQDM BigGAN Avg.
Midj. 88.1 87.0 86.9 84.7 88.3 88.3 87.3 86.3 87.1
SDV1.4 88.3 89.9 89.6 83.2 86.7 62.6 83.8 88.3 84.0
SDV1.5 89.4 89.7 90.4 84.4 89.8 87.1 87.2 88.7 88.3
ADM 83.0 82.5 83.5 82.4 83.6 83.7 84.5 82.6 83.2
Glide 85.7 83.9 84.1 85.4 87.4 87.4 87.8 83.8 85.7
Wukong 81.2 75.5 76.7 84.5 89.5 88.6 87.8 77.8 82.7
VQDM 83.0 83.4 84.1 83.6 84.5 84.7 85.5 83.0 84.0
BigGAN 88.1 87.8 88.5 84.2 88.3 88.7 88.7 87.7 87.7
Avg. 85.8 85.0 85.5 84.0 87.3 83.9 86.5 84.8 85.3
Table 2: Cross-generator evaluation of $k$-NN on the GenImage validation set for fake image detection. We report the accuracy of the selected $k$-NN configuration. For each generator, we use the support set from one model and test it on the validation images of all the generators.
Support set from Tested on
Midj. SDV1.4 SDV1.5 ADM Glide Wukong VQDM BigGAN Avg.
Midj. 91.4 89.1 89.3 88.0 91.5 91.6 89.9 88.6 89.9
SDV1.4 89.9 90.9 90.7 85.4 88.2 64.1 84.7 89.4 85.4
SDV1.5 91.7 91.7 91.4 86.7 91.9 89.9 89.4 90.9 90.4
ADM 86.2 85.5 85.6 86.7 86.9 86.5 86.5 85.1 86.1
Glide 89.0 86.7 86.8 89.6 91.1 90.6 90.0 87.2 88.9
Wukong 85.3 78.7 79.5 88.2 93.0 92.7 91.2 81.6 86.3
VQDM 86.7 86.6 86.5 87.9 88.6 88.1 87.9 85.8 87.3
BigGAN 90.8 90.3 89.9 87.4 90.9 91.6 90.9 89.9 90.2
Avg. 88.9 87.4 87.5 87.5 90.3 86.9 88.8 87.3 88.1
Table 3: Cross-generator evaluation of $k$-NN on the GenImage test set for fake image detection. We report the accuracy of the selected $k$-NN configuration. For each generator, we use the support set from one model and test it on the images of all the generators.
Methods Tested on Avg.
Midjourney SDV1.4 SDV1.5 ADM GLIDE Wukong VQDM BigGAN
CNNSpot [46] 58.2 70.3 70.2 57.0 57.1 67.7 56.7 56.6 61.7
Spec [56] 56.7 72.4 72.3 57.9 65.4 70.3 61.7 64.3 65.1
F3-Net [33] 55.1 73.1 73.1 66.5 57.8 72.3 62.1 56.5 64.6
GramNet [26] 58.1 72.8 72.7 58.7 65.3 71.3 57.8 61.2 64.7
DIRE [49] 65.0 73.7 73.7 61.9 69.1 74.3 63.4 56.7 67.2
LaRE2 [29] 66.4 87.3 87.1 66.7 81.3 85.5 84.4 74.0 79.1
LATTE [43] 71.3 79.3 81.8 82.2 92.8 82.0 82.9 87.8 82.5
FRIDA (Ours) 88.9 87.5 87.5 87.5 90.3 86.9 88.9 87.4 88.1
Table 4: Fake image detection comparison with state-of-the-art approaches on GenImage test set. The reported metric is the average percentage accuracy calculated from eight distinct models, each trained on the data from a different generator, and tested on the specified test set. † indicates that the results are reproduced by [29]. Best in bold, second best underlined.
As detailed in Table 4, our proposed approach demonstrates a significant leap in performance over current state-of-the-art methods on the GenImage test set. The experimental setup is designed to rigorously test generalization: eight models are trained on limited data (2000 samples) from a single generator each, but evaluated on images from all generators. In this challenging scenario, our method is the one that better generalizes across these unseen data distributions, establishing a new state-of-the-art performance by a margin of nearly six percentage points.
Beyond its accuracy, the primary advantage of our approach lies in its remarkable data and training efficiency. The strong results are achieved without a conventional training process; we use a pre–trained SDM and the $k$-NN framework merely requires storing the support set. The proposed approach maintains high performance even with a drastically smaller support set. When the sample size is reduced by $90\%$, the average cross-generator accuracy on the test set drops by only $1.6\%$ (from $88.1\%$ to $86.5\%$), a result that still outperforms the previous state-of-the-art (Table 4). This characteristic is crucial for practical deployment. In a field where new generative models are constantly emerging, our method provides a scalable and effective solution that does not require cost-prohibitive data generation and retraining cycles. Using an NVIDIA RTX 4090, SDM feature extraction required an average of 2.1 seconds, while $k$-NN classification took 0.0003 seconds. In contrast, on a CPU (Intel Core i7-9800X @ 3.80 GHz), feature extraction averaged 12.9 seconds and classification 0.002 seconds.
5.3 Source Model Attribution
We compare MLP and $k$-NN classifiers on their ability to identify an image’s source generator, following the experimental procedure described in Section 4.3.
$k$-NN Classifier.
A portion of the results from the hyperparameter selection on the validation set obtained using the correlation distance is presented in Table 5.
$k$ Support sizes 18 90 180 900 1800 3600 9000 1 25.2% 33.5% 36.4% 44.6% 48.1% 51.6% 55.8% 9 28.7% 31.3% 44.09% 48.7% 52.7% 57.7% 19 27.2% 42.5% 48.3% 52.1% 57.1% 35 39.5% 46.5% 50.9% 56.1% 101 39.7% 46.3% 52.0%
Table 5: Effect of $k$ and support set size on source attribution accuracy. Validation accuracies for the nine-class attribution task (eight generators plus real images) computed across different values of $k$ and support set sizes.
These results proved that the $k$-NN classifier is inadequate for effective source image attribution. Despite optimization (using a support set of 9,000 and k=9), its peak accuracy reached a mere $57.7\%$, a performance level far too low for practical application.
Neural Network Classifier.
We train the three MLP classifiers as described in Section 4.3; in Table 6 we report the average results of the MLPs on the validation set across ten training runs.
Model Avg. Accuracy Std.
MLP-640 84.87% 0.143
MLP-320 84.71% 0.205
MLP-640-320 84.78% 0.127
Table 6: Source attribution performance of MLP models. Average accuracy and standard deviation over ten runs on the nine-class attribution task (eight generators plus real images) on the validation set.
While the average accuracies of the three models on the source attribution task are comparable, the MLP-640, with $84.87\%$ of average accuracy, slightly outperforms the others and allows for an increase in performance of about 27 percentage points over the $k$-NN classifier. The MLP-640, when tested on the GenImage test set, achieves an accuracy of $84.36\%$. This result is consistent with the validation set performance, indicating a high degree of generalization.
Figure 3: Confusion Matrix for source image attribution. The MLP-640 is evaluated on the GenImage test set.
As shown in Figure 3, the classifier generally recognizes source generators with high accuracy; however, it struggles to distinguish models that share the same architecture. In particular, it frequently confuses SDM v1.4 with SDM v1.5, and, in some cases, the Wukong model — which is based on SDM — is also mistaken for them.
Notably, the huge performance increase of the MLP over the $k$-NN classifier on the validation set confirms that the features extracted by the diffusion U-Net contain generator-specific patterns. The MLP is capable of learning these distinguishing characteristics, whereas the $k$-NN approach, which relies solely on feature distances, cannot recognize these subtle patterns. Although $k$-NN proves effective for the broader task of distinguishing real from synthetic content, it is insufficient for the more nuanced task of source attribution. This suggests that while the feature distance between real and synthetic images is large, the distances between features from different generators are smaller. The key differentiators are not the absolute distances but rather specific, patterns within the features that the MLP can successfully identify.
To investigate this hypothesis, we employ the SHAP algorithm to interpret the decisions of the MLP-640 classifier. We utilize the Gradient Explainer with a background dataset of 200 samples and 1000 test samples. For each class, we identify the top 10 most informative features. In Figure 4 we report the percentage of the top 10 features shared between each pair of generators.
Figure 4: Shared informative features across generators. We use SHAP with the Gradient Explainer to interpret the decisions of the MLP-640 classifier, employing 200 background and 1000 test samples. The plot reports, for each pair of generators, the percentage of overlap among their top 10 most informative features.
[Real]Refer to caption [ADM]Refer to caption [BigGAN]Refer to caption
[Glide]Refer to caption [Midjourney]Refer to caption [SDV1.4]Refer to caption
[SDV1.5]Refer to caption [VQDM]Refer to caption [Wukong]Refer to caption
Figure 5: Impact of the top-10 features on model decisions by SHAP analysis. Visualization of the ten most influential features identified by the SHAP analysis for each of the eight diffusion generators and the real image class. The figure highlights how different feature subsets contribute to the network’s decision process across generators, revealing both shared and model-specific attribution patterns.
The two SDMs exhibit a $50\%$ overlap in their top 10 most important features. Furthermore, these shared features influence the model’s output identically (see Figure 5 and Figure 5), indicating that the images generated by these two models lead to a latent representation with common characteristics. As a result, the images generated by the two models are nearly indistinguishable to MLP. In contrast, the classifier successfully discriminates real and BigGAN images from the SDMs by exploiting divergent feature behavior. Although these models share $30\%$ of their top 10 features with the SDMs, the impact of these common features is markedly different (Figure 5 and Figure 5). Finally, the Wukong model presents an interesting paradox. Despite occasional misclassifications with the SDM-generated images, the SHAP analysis reveals zero feature overlap within the top 10 most influential features (Figure 5). This indicates that the classifier’s errors are not driven by the same high-impact features that define the SDM class. Instead, the confusion likely arises because a different set of features in Wukong images produces a combined effect that coincidentally mimics the characteristics of an SDM, leading the classifier to an incorrect conclusion.
6 Conclusions
In this work, we presented FRIDA, a training-free framework that repurposes internal activations of pre-trained diffusion models for deepfake detection and source generator attribution. FRIDA operates entirely in the diffusion feature space, leveraging the latent representation of a Stable Diffusion U-Net as discriminative descriptors of image authenticity. A simple k-nearest-neighbor classifier applied to these features achieves state-of-the-art detection performance and generalizes effectively across unseen generators, while a compact neural model trained on the same representations enables accurate source attribution. Our analysis shows that diffusion features, although optimized for image generation, are surprisingly effective for fake image detection and source model attribution. This finding suggests that diffusion features can serve as a universal and interpretable basis for synthetic image forensics, bridging generative modeling and authenticity analysis.

### v2

@@
 Abstract
-The rapid progress of generative diffusion models has enabled the creation of synthetic images that are increasingly difficult to distinguish from real ones, raising concerns about authenticity, copyright, and misinformation.
-Existing supervised detectors often struggle to generalize across unseen generators, requiring extensive labeled data and frequent retraining.
-We introduce FRIDA (Fake-image Recognition and source Identification via Diffusion-features Analysis), a lightweight framework that leverages internal activations from a pre-trained diffusion model for deepfake detection and source generator attribution.
-A k-nearest-neighbor classifier applied to diffusion features achieves state-of-the-art cross-generator performance without fine-tuning, while a compact neural model enables accurate source attribution.
-These results show that diffusion representations inherently encode generator-specific patterns, providing a simple and interpretable foundation for synthetic image forensics.
-Index Terms —
-Overview of our framework for deepfake classification and attribution.
+The rapid rise of generative models has yielded synthetic images of striking realism, blurring the line between real and fake content.
+As novel models proliferate, detectors must go beyond mere fake identification to robustly generalise across unseen generators and synthetic content.
+We introduce FRIDA (Fake image Recognition and source Identification via Diffusion features Analysis), a lightweight, data-efficient framework that uses features from a pre-trained Stable Diffusion Model to detect and attribute AI-generated images.
+Through an in-depth analysis of how data from different generators are encoded across diffusion U-Net layers, we propose a method that (i) detects synthetic images using a training-free $k$-Nearest Neighbour approach and (ii) performs source model attribution via a compact neural classifier.
+On the GenImage benchmark, FRIDA achieves state-of-the-art cross-generator detection with limited data while maintaining robust source model attribution capabilities.
+These results establish diffusion features as a reliable framework for AI-generated image forensics.
+Overview of our framework for fake image classification and source model attribution.
 Given an input image, our Image Prototyping module extracts image prototype from the U-Net of a Stable Diffusion Model.
-This compact representation is then used for two downstream tasks: (i) Deepfake classification via a k-NN classifier to determine whether the image is real or fake, and (ii) Deepfake attribution via an MLP to identify the generative model (e.g., GLIDE, Stable Diffusion, BigGAN) responsible for the fake.
+This compact representation is then used for two downstream tasks: (i) fake image classification via a $k$-NN classifier to determine whether the image is real or fake, and (ii) source model attribution via an MLP to identify the generative model (e.g., GLIDE, Stable Diffusion, BigGAN) responsible for the fake.
 1 Introduction
-The outstanding capabilities of recent generative models to synthesize media with exceptional fidelity and variety [1, 16, 4, 36] have accelerated the proliferation of AI‑generated content across domains, including medical imaging [23, 47, 42], education [53, 51], marketing [20, 12], and robotics [50].
-This rapid progress has been largely driven by diffusion-based generators — including Stable Diffusion Model (SDM) [36], DALL-E  [35], and Imagen [37] — which produce highly realistic and semantically coherent visuals from simple text prompts.
+The outstanding capabilities of recent generative models to synthesize media with exceptional fidelity and variety [midjourney, nanobanana, betker2023improving, Rombach_2022_CVPR] have accelerated the proliferation of AI‑generated content across domains, including medical imaging [KAZEROUNI2023102846, WANG2025102593, 10937272], education [xia2024scoping, wu2025systematic], marketing [grewal2025generative, cillo2025generative], and robotics [wolf2025diffusion].
+This rapid progress has been largely driven by diffusion-based generators—including Stable Diffusion Model (SDM) [Rombach_2022_CVPR], DALL-E  [ramesh2021zero], and Imagen [saharia2022photorealistic]—which produce highly realistic and semantically coherent visuals from simple text prompts.
 This increasing accessibility of advanced image generation tools through user-friendly interfaces poses considerable risks, including the creation of biased content and the violation of intellectual property rights.
-To mitigate these risks, robust detection and attribution methods are required to reliably identify synthetic media and determine their source models [39, 13, 29].
-Early fake detection methods relied on supervised learning pipelines [52, 59, 29], which depend on vast labelled datasets and intensive computation.
-Approaches for source generator attribution pushed these requirements even further, as they must learn generator-specific fingerprints [54, 8, 46].
+To mitigate these risks, robust detection and attribution methods are required to reliably identify synthetic media and determine their source models [sha2023fake, luo2024lare, cioni2024clip].
+Early fake detection methods relied on supervised learning pipelines [wu2025explainable, zhu2023genimage, luo2024lare], which depend on vast labelled datasets and intensive computation.
+Approaches for source model attribution pushed these requirements even further, as they must learn generator-specific fingerprints [yu2019attributing, bui2022repmix, wang2020cnn].
 However, as generative models evolve rapidly, the continual need for data collection and retraining makes classifier-based detectors hard to scale, underscoring the need for lightweight, data-efficient methods that generalize across both known and emerging generators.
-An opportunity to move beyond task-specific classifiers has emerged with the rise of large-scale foundation models based on Vision Transformers (ViTs) [19], e.g., CLIP [34] and DINO [9].
-Pre-trained on massive datasets, these models learn general, transferable representations that have shown remarkable effectiveness in downstream tasks [24, 45] — including deepfake detection [31, 40, 44] — when their extracted features are used for classification.
+An opportunity to move beyond task-specific classifiers has emerged with the rise of large-scale foundation models based on Vision Transformers (ViTs) [dosovitskiy2021an], e.g., CLIP [radford2021learning] and DINO [caron2021emerging].
+Pre-trained on massive datasets, these models learn general, transferable representations that have shown remarkable effectiveness in downstream tasks [kerssies2025your, wang2024enhancing]—including deepfake detection [nguyen2024exploring, smeu2025declip, wang2025fsfm]—when their extracted features are used for classification.
 Similarly, diffusion-based generative models, though designed for image generation, have proven to be remarkably effective feature encoders.
-Recent studies have shown that pre-trained SDMs provide rich semantic information about image content [5, 6], enabling strong performance in applications such as semantic segmentation [3, 15], and other vision-related problems — substantially reducing the computational cost of training large models from scratch.
+Recent studies have shown that pre-trained SDMs provide rich semantic information about image content [BONECHI20245244, BONECHI2025128846], enabling strong performance in applications such as semantic segmentation [baranchuk2021label, corradini2024freeseg], and other vision-related problems—substantially reducing the computational cost of training large models from scratch.
 At the same time, diffusion models have been increasingly used for fake detection.
-Some methods identify diffusion-generated images through reconstruction error [49], while others analyse latent trajectories via diffusion inversion to enhance generalisation across unseen generators [10, 43].
+Some methods identify diffusion-generated images through reconstruction error [wang2023dire], while others analyse latent trajectories via diffusion inversion to enhance generalisation across unseen generators [cazenavette2024fakeinversion, vasilcoiu2025latte].
 Based on the principle that images from different sources exhibit distinct distributions, Zhong et al.
-[57] employ a SDM as a denoising tool to identify fake images by detecting the resulting artifacts in their reconstructed versions.
+[zhong2025beyond] employ a SDM as a denoising tool to identify fake images by detecting the resulting artifacts in their reconstructed versions.
 However, these approaches typically require repeated diffusion or inversion steps, making them computationally expensive and less practical for large-scale or real-time detection scenarios.
-Inspired by these recent advances, we explore whether the internal features of a pre-trained diffusion-based generative model can effectively separate real from synthetic images and, beyond detection, enable attribution to the source generator.
-To this end, we introduce a lightweight and data-efficient framework that operates entirely within the latent feature space of SDMs at the last inference step.
-Our approach detects deepfakes by applying a training-free $k$-Nearest Neighbours ($k$-NN) classifier to latent representations extracted from the SDM.
-This approach achieves state-of-the-art performance on the GenImage benchmark  [59] for fake image detection.
+Inspired by these recent advances, we explore whether the internal features of a pre-trained diffusion-based generative model can effectively separate real from synthetic images drawn by various generators and, beyond detection, enable attribution to the source generator.
+To this end, we propose FRIDA, a data-efficient framework for detecting fake images.
+The core of our method is to repurpose a pre-trained SDM (v1.5) at the last inference step as a feature extractor.
+Then, these features are employed by a simple $k$-Nearest Neighbours ($k$-NN) classifier, showing state-of-the-art results on the GenImage benchmark [zhu2023genimage].
 This strategy eliminates the need for costly inversion or fine-tuning and shows strong generalisation to unseen generators.
-We further extend our method to source-model attribution, showing that a lightweight Multi-Layer Perceptron (MLP) trained on the SDM latent features can accurately identify the generator responsible for a synthetic image.
+We further extend our method to source model attribution, showing that a lightweight Multi-Layer Perceptron (MLP) trained on the SDM latent features can accurately identify the generator responsible for a synthetic image.
 The main contributions of our work can be summarized as follows:
   1. 1.
-    We show that latent features extracted from specific layers of a pre-trained SDM are highly discriminative for both fake image detection and source model attribution.
+    We show that latent features extracted from specific layers of a pre-trained SDM are highly discriminative for fake image detection.
+Furthermore, we benchmark the efficacy of SDM as a feature extractor against prominent ViT-based backbones (CLIP and DINO), establishing that SDM-based representations achieve superior detection performance.
   2. 2.
-    We benchmark our fake image detection approach on data generated by eight different models, achieving state-of-the-art performance with a novel, lightweight $k$-NN-based framework that is entirely training-free.
+    We evaluate our fake image detection approach on data generated by eight different models, achieving state-of-the-art performance on GenImage dataset with a novel $k$-NN-based framework that is entirely training-free.
+We further demonstrate the robustness of our method by testing on a challenging unseen subset comprising state-of-the-art generators, including Flux and SDv3.5.
 Our method requires only a small support set, generalises effectively to unseen generators, and adapts to new data without any fine-tuning.
   3. 3.
+    We evaluate the robustness of FRIDA against varying levels of image perturbations, including Gaussian noise and JPEG compression, to simulate real-world distortions.
+  4. 4.
     A lightweight neural classifier trained on the SDM latent representation proves to be highly effective in source model attribution, suggesting the presence of generator-specific characteristics in the features.
-We use SHAP (SHapley Additive exPlanations) [28] to further investigate the capability of the SDM to encode these model-specific signatures.
-The remainder of this paper is organized as follows:  Section 2 reviews relevant literature on synthetic image detection and attribution.
-Section 3 describes the GenImage dataset and the procedure used to extract the latent representation of each image using the SDM.
-Section 4 describes the experimental setup, while  Section 5 presents and discusses the obtained results.
-Finally,  Section 6 concludes the paper with a summary of our findings and outlines future research directions.
+We use SHAP (SHapley Additive exPlanations) [lundberg2017unified] to further investigate the capability of the SDM to encode these model-specific signatures.
+The remainder of this paper is organized as follows:  Section˜2 reviews relevant literature on synthetic image detection and attribution.
+Section˜3 presents the benchmark dataset (GenImage) alongside an auxiliary dataset of recent generators, and outlines the methodology for extracting per-image latent representations using the SDM.
+Section˜4 describes the experimental setup, while  Section˜5 presents and discusses the obtained results.
+Finally,  Section˜6 concludes the paper with a summary of our findings and outlines future research directions.
 2 Related Work
@@
 Zhang et al.
-[56] introduced AutoGAN, a frequency-domain simulator of upsampling artifacts, and trained a spectrum-based classifier to detect periodic Fourier patterns without access to the generator.
+[zhang2019detecting] introduced AutoGAN, a frequency-domain simulator of upsampling artifacts, and trained a spectrum-based classifier to detect periodic Fourier patterns without access to the generator.
 Similarly, Wang et al.
-[46] showed that CNN-generated images share common low-level artifacts across architectures and datasets.
-Their ResNet-based detector, trained on ProGAN, generalized remarkably well to unseen generators such as StyleGAN and BigGAN, revealing universal CNN fingerprints that persist across models.
-Gram-Net [26] models global texture correlations via multi-layer Gram matrices, improving robustness to compression and noise and enabling better cross-GAN generalization by leveraging long-range texture cues.
+[wang2020cnn] showed that CNN-generated images share common low-level artifacts across architectures and datasets.
+Their ResNet-based detector, trained on images from ProGAN, generalized remarkably well to unseen generators such as StyleGAN and BigGAN, revealing universal CNN fingerprints that persist across models.
+Gram-Net [liu2020global] models global texture correlations via multi-layer Gram matrices, improving robustness to compression and noise and enabling better cross-GAN generalization by leveraging long-range texture cues.
 Qian et al.
-[33] advanced this line of work with F3-Net, a two-branch frequency-aware CNN classifier combining frequency decomposition and local statistics via cross-attention, outperforming spatial-domain detectors on a deepfake faces dataset even under strong compression.
+[qian2020thinking] advanced this line of work with F3-Net, a two-branch frequency-aware CNN classifier combining frequency decomposition and local statistics via cross-attention, outperforming spatial-domain detectors on a deepfake faces dataset even under strong compression.
 Diffusion-based Detection and Reconstruction Methods.
-Following the success of diffusion-based models, several works have focused on detecting diffusion-generated images.
-DE-FAKE [38] is a hybrid two-branch architecture that exploits image-text consistency between captions and visual content to detect synthetic images.
-DIRE [49] detects diffusion-generated images by inverting an input into the diffusion latent space and then reconstructing it through the full reverse denoising trajectory; the discrepancy between the original and reconstructed signals is used for detection.
-Similarly, LaRE2 [29] estimates a Latent Reconstruction Error using only a single-step reconstruction in latent space, using a module that aligns and refines features across spatial and channel dimensions.
-A related line of work [48] also leverages a pre-trained diffusion model, extracting multi-timestep responses under the hypothesis that synthetic images, which lie outside the natural image manifold, exhibit distinctive denoising behavior.
-ESIDE [52] incorporates frequency perturbations into diffusion inversion, training an ensemble of CLIP-based classifiers on noised representations to improve robustness and interpretability.
-More recently, LATTE [43] introduces a Latent Trajectory Embedding framework that explicitly models the temporal evolution of latent representations across denoising steps.
-By aggregating multi-timestep latent features through joint visual–latent refinement, LATTE captures dynamic generation cues beyond reconstruction error, achieving strong cross-generator and cross-dataset generalization.
-An alternative approach [57] treats detection as an anomaly detection task.
-They learn the low-level feature distribution of real images by training an extractor to spot pixel-level differences between original images and their denoised counterparts, effectively identifying generated content that falls outside this learned distribution.
+Following the success of diffusion-based models, several works have focused on detecting images generated by such models.
+DE-FAKE [defake2023] is a hybrid two-branch architecture that exploits image-text consistency between captions and visual content to detect synthetic images.
+DIRE [wang2023dire] detects diffusion-generated images by inverting an input into the diffusion latent space and then reconstructing it through the full reverse denoising trajectory; the discrepancy between the original and reconstructed signals is used for detection.
+Similarly, LaRE2 [luo2024lare] estimates a Latent Reconstruction Error using only a single-step reconstruction in latent space, using a module that aligns and refines features across spatial and channel dimensions.
+A related line of work [diffimplicit2025] also leverages a pre-trained diffusion model, extracting multi-timestep responses under the hypothesis that synthetic images, which lie outside the natural image manifold, exhibit distinctive denoising behavior.
+ESIDE [wu2025explainable] incorporates frequency perturbations into diffusion inversion, training an ensemble of CLIP-based classifiers on noised representations to improve robustness and interpretability.
+More recently, LATTE [vasilcoiu2025latte] introduces a Latent Trajectory Embedding framework that explicitly models the temporal evolution of latent representations across denoising steps.
+By aggregating multi-timestep latent features through joint visual-latent refinement, LATTE captures dynamic generation cues beyond reconstruction error, achieving strong cross-generator and cross-dataset generalization.
+An alternative approach [zhong2025beyond] treats detection as an anomaly detection task.
+It learns the low-level feature distribution of real images by training an extractor to spot pixel-level differences between original images and their denoised counterparts, effectively identifying generated content that falls outside this learned distribution.
 Semantic and Open-Set Attribution Frameworks.
-Recent research explores semantic alignment and open-set detection to generalize across diverse generators.
+To generalize across diverse generators, recent research explores semantic alignment, which disentangles generator-specific artifacts from image semantics, and open-set detection, which extends attribution to generators unseen at training time.
 Zhu et al.
-[58] propose MAID, a framework-agnostic attribution method that extracts Diffusion Model Activations (DMA) by treating pre-trained diffusion models as denoising autoencoders.
+[zhu2025maid] propose MAID, a framework-agnostic attribution method that extracts diffusion model activations by treating pre-trained diffusion models as denoising autoencoders.
 These activations encode model-specific patterns without requiring white-box access or prompts, supporting both detection and attribution.
-SemGIR [55] employs semantic-guided image regeneration: a candidate image is captioned, regenerated via text-to-image synthesis, and compared with its reconstruction using CLIP-based encoders.
+SemGIR [semgir2024] employs semantic-guided image regeneration: a candidate image is captioned, regenerated via text-to-image synthesis, and compared with its reconstruction using CLIP-based encoders.
 This forces the detector to focus on generator-specific artifacts rather than prompt semantics and achieves strong cross-generator generalization.
 Cioni et al.
-[14] move beyond frequency-based fingerprints by leveraging intermediate representations of large ViT-based models such as CLIP and DINOv2.
-Their open-set attribution framework combines linear probing and k-nearest neighbors with confidence- or distance-based rejection to identify images from unseen generators, providing a unified and retraining-free solution across GAN and diffusion sources.
-3 Materials and Methods
-3.1 GenImage Dataset
-GenImage [59] is a large-scale benchmark designed to detect generated imagery, composed of both synthetic and real images sourced from ImageNet [17].
-The dataset contains roughly 1.35 million synthetic images produced by eight distinct generative models: BigGAN [7], GLIDE [32], VQDM [21], SDM (v1.4 and v1.5)[36], ADM[18], Midjourney [1], and Wukong [2].
-Both the real and synthetic images cover the 1,000 ImageNet classes, with the synthetic portion providing approximately 1,350 images per class (1,300 for training and 50 for testing).
+[clipfeat2024] move beyond frequency-based fingerprints by leveraging intermediate representations of large ViT-based models such as CLIP and DINO.
+Their open-set attribution framework combines linear probing and $k$-nearest neighbors with confidence-based or distance-based rejection to identify images from unseen generators, providing a unified and retraining-free solution across GAN and diffusion sources.
+3 Dataset Overview and Image Prototyping
+3.1 Datasets
+GenImage Dataset.
+GenImage [zhu2023genimage] is a large-scale benchmark designed to detect generated imagery, composed of both synthetic and real images sourced from ImageNet [deng2009imagenet].
+The dataset contains roughly 1.35 million synthetic images produced by eight distinct generative models: BigGAN [brock2018large], GLIDE [nichol2022glide], VQDM [gu2022vector], SDM (v1.4 and v1.5) [Rombach_2022_CVPR], ADM [dhariwal2021diffusion], Midjourney [midjourney], and Wukong [wukong].
+Both the real and synthetic images cover the 1,000 ImageNet classes, with the synthetic portion providing 1,350 images per class (1,300 for training and 50 for testing).
 The generated images were created using simple text prompts following the template “photo of [class]” and have resolutions ranging from $128\times 128$ to $1024\times 1024$, depending on the source model.
@@
 Finally, the original GenImage test set was used as a held-out evaluation set.
+Other Datasets.
+To evaluate generalization beyond GenImage, we assembled an alternative, out-of-distribution (OOD) test set by extracting 500 synthetic images per generator and 500 real images per dataset from different sources.
+In particular, fake images are sampled from the Qwen-Image-Self-Generated-Dataset [wu2025qwenimagetechnicalreport] and from five generators of the OpenFake dataset [livernoche2025openfake] (i.e., Flux 1.1 Pro, Flux 1.0 Dev, GPT Image 1, Ideogram 3.0, and SDv3.5).
+The real images are sampled from LAION 400M [schuhmann2021laion] and Pascal VOC [Everingham15].
+3.2 The GenImage Evaluation Protocol
+The cross-generator evaluation protocol used in this work follows the one in  [zhu2023genimage].
+To assess generalization, a binary real-vs-fake classifier is trained on one generator-specific subset and then tested—without any adaptation—on evaluation data from both the same generator (seen data distribution) and from the other seven generators (unseen data distributions).
 3.3 Image Prototype Extraction
-We exploit a pre-trained SDM v1.5111https://github.com/hkproj/pytorch-stable-diffusion and we use the following procedure for extracting internal features prototype from a specific layer (see Figure 1).
+We adopt a pre-trained SDMv1.5111https://github.com/hkproj/pytorch-stable-diffusion and we use the following procedure for extracting internal features prototype from a specific layer (see Figure˜1).
 Figure 1: Prototype extraction from Stable Diffusion U-Net. In this example, we extract and average the features from the first decoder layer at $16\times 16$ resolution.
@@
 This section details the experimental setup used in our study.
-First, we compare features from different layers of the diffusion U-Net to identify which layer is most effective at distinguishing real from synthetic content.
-We then describe the experimental setups for our two primary tasks: fake image detection (Section 4.2) and source generator attribution (Section 4.3).
+First, we compare features from different layers of the diffusion U-Net to identify which layer is most effective at distinguishing real from synthetic content (Section˜4.1).
+We then describe the experimental setups for our two primary tasks: fake image detection (Section˜4.2) and source model attribution (Section˜4.3).
+Finally, in Section˜4.4, we present additional studies using different backbones, out-of-distribution datasets and various image perturbations.
 4.1 U-Net Layer Selection by Linear Probing
-To identify which layers provide the most effective features for deepfake detection, we extract latent representations from the encoder, bottleneck, and decoder layers of the U-Net at multiple spatial resolutions ($64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$).
-Following the GenImage protocol [59], we train eight classifiers, one for each generator in the dataset, and evaluate each classifier on images from all other generators.
-In particular, we adopt a linear probing setup — commonly used to evaluate the quality and separability of learned representations in self-supervised and foundation models  [11, 22, 34] — by training a linear classifier with sigmoid activation on the latent prototypes from each layer.
-Each classifier is trained on the training subset (see Section 3.1) with real and synthetic images from a given generator.
-All models are trained with the AdamW optimizer [27] using a learning rate of $3\times 10^{-4}$.
+To identify which layers provide the most effective features for fake image detection, we extract latent representations from the encoder, bottleneck, and decoder layers of the U-Net at multiple spatial resolutions ($64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$).
+We adopt a linear probing setup—commonly used to evaluate the quality and separability of learned representations in self-supervised and foundation models  [chen2020simple, he2020momentum, radford2021learning]—by training a linear classifier with sigmoid activation on the latent prototypes from each layer.
+Each classifier is trained on the training subset (see Section˜3.1) with real and synthetic images from a given generator.
+All models are trained with the AdamW optimizer [loshchilov2018decoupled] using a learning rate of $3\times 10^{-4}$.
 Training is early stopped based on validation accuracy, with a patience of 15 epochs.
-To assess generalization, we adopt a cross-generator evaluation: following common practice [59], each classifier trained on a specific generator is evaluated on the fake detection task using its own validation set (seen data distribution) as well as the validation sets of the seven remaining generators (unseen data distributions).
-The U-Net layer whose features yield the highest average cross-generator accuracy is then selected and used for all subsequent experiments.
+To assess generalization, we perform the cross-generator evaluation described in Section˜3.2 using the eight validation sets.
+The U-Net layer whose features yield the highest average cross-generator validation accuracy is then selected and used for all subsequent experiments.
 4.2 Fake Image Detection
@@
 To this aim, we evaluate two approaches: a simple neural network (MLP) and a distance-based classifier ($k$-NN).
-In both cases, we follow the cross-generator evaluation protocol described in Section 4.1, using the selected U-Net layer.
+In both cases, we extract the features from the selected U-Net layer and follow the cross-generator evaluation protocol described in Section˜3.2 to select the best model configurations based on the validation accuracy.
+The best MLP and $k$-NN models are then tested on the GenImage test set.
 MLP for Fake Image Detection.
@@
 All the models have two output neurons with a softmax activation function and are trained using binary cross-entropy loss with AdamW optimizer and a learning rate of $3\times 10^{-4}$.
-The validation set accuracy is used to early stop the training process with a patience of 15 epochs.
+The validation accuracy is used to early stop the training process with a patience of 15 epochs.
 $k$-NN for Fake Image Detection.
 […]
-The best MLP and $k$-NN models are selected following the cross-generator protocol described in Section 4.1, based on validation accuracy, and then tested on the GenImage test set.
 4.3 Source Model Attribution
-Our second objective is source generator attribution, a multi-class classification task to identify the source model of an image.
+Our second objective is image attribution, a multi-class classification task to identify the source model of an image.
 As for fake image detection, we compare the MLP and $k$-NN approaches, adapting the models and the evaluation protocol for multi-class classification.
-More specifically, we employ a single classifier with nine output classes (eight generators and real images).
+More specifically, we employ a single classifier with nine output classes (eight generators + real images).
 MLP for Source Model Attribution.
-We evaluate the same MLP configurations used for fake-image detection (see Section 4.2), but replace the output layer with a 9-way softmax classifier and train with cross-entropy loss.
-All MLPs are trained on synthetic images from the training subsets of all generators, while real images are sampled from the training subset of the Midjourney dataset.
+We evaluate the same MLP configurations used for fake-image detection (see Section˜4.2), but we replace the output layer with a 9-way softmax classifier and train with cross-entropy loss.
+All MLPs are trained on synthetic images from the training subsets of all generators, while the real images are sampled from the training subset to maintain the class balance.
 All models are trained using the AdamW optimizer with a learning rate of $3\times 10^{-4}$, and the validation accuracy is used to early stop the training process using a patience of 15 epochs.
 $k$-NN for Source Model Attribution.
-[…]
-The most effective approach is then evaluated on the GenImage test set for the attribution task.
+We perform the hyperparameter selection for $k$-NN using the same grid employed in Section˜4.2, changing the sizes of the support set $S$ (ranging from 18 to 9,000) [^4] Sizes(S)=Sizes(S)={18, 45, 90, 180, 360, 450, 900, 1350, 1800, 2700, 3600, 4500, 9000} to comprise an equal number of images for all nine classes.
+4.4 Extended Experimental Analysis
+Benchmarking Against ViT Backbones.
+We assess the advantage of using SDM features in our method by conducting a comparative analysis against established ViT-based feature extractors, specifically CLIP and DINO.
+For both, we extract the features from their last layer (widely used for their robust performance across various downstream tasks) and we evaluate the cross-generator performance under the same setup used for the SDM.
+Robustness to Image Corruptions.
+We evaluate the robustness of FRIDA by introducing various synthetic distortions to the test set.
+Specifically, we assess the cross-generator performance across multiple degradation levels, including JPEG compression (10, 30, 50, 70, 90), gaussian noise ($\sigma$=0.05, 0.1, 0.15, 0.2, 0.25), and gaussian blur ($\sigma$=1, 2, 3, 4, 5).
+Out-of-Distribution Robustness.
+We further evaluate the generalization capabilities of our approach on OOD datasets.
+The support set for $k$-NN is a random subset of the GenImage dataset encompassing all available generators and their corresponding real images.
+During inference, we evaluate the model against a series of unseen OOD generators.
+To isolate the domain shift introduced by the synthetic data, the real images in the test set remain constant across all evaluations, comprising 500 samples balanced equally between Pascal VOC ($N=250$) and LAION-400M ($N=250$).
+For each test iteration, these are paired with 500 synthetic images generated by a single OOD model, thereby maintaining a strictly balanced binary classification setup.
 5 Results
-This section presents our experimental results, starting with the analysis of features extracted from the diffusion U-Net (Section 5.1), and then covering fake image detection (Section 5.2) and source model attribution (Section 5.3).
+This section presents our experimental results, starting with the U-Net layer selection for feature extraction (Section˜5.1), and then covering fake image detection (Section˜5.2) and source model attribution (Section˜5.3).
+Finally, in Section˜5.4, we present the results of the extended analysis using different backbones, OOD datasets and various image perturbations.
 5.1 U-Net Layer Selection by Linear Probing
-Following the experimental setup in Section 4.1, we test image prototypes from all the layers of the diffusion U-Net using a linear probing approach to determine (i) if the latent space representation can be used to distinguish between real and synthetic data, and (ii) which layer produces the most informative representation for this task.
-Figure 2 plots the average cross-generator validation accuracy for features extracted from different levels of the diffusion U-Net.
+Following the experimental setup in Section˜4.1, we test image prototypes from all the layers of the diffusion U-Net using a linear probing approach to determine (i) if the latent space representation can be used to distinguish between real and synthetic data, and (ii) which layer produces the most informative representation for this task.
+Figure˜2 plots the average cross-generator validation accuracy for features extracted from different levels of the diffusion U-Net.
 Figure 2: U-Net Layer Selection by Linear Probing.
@@
 The best accuracy is achieved by features from the first layer of the decoder at $16\times 16$ resolution.
-The results of our feature-probing analysis show that the latent features extracted from the diffusion U-Net can be highly discriminative for distinguishing real from synthetic images.
+The results of our feature probing analysis show that the latent features extracted from the diffusion U-Net can be highly discriminative for distinguishing real from synthetic images.
 In particular, we identify the first layer of the decoder at $16\times 16$ resolution (Decoder 16_0) as the source of the most informative features, yielding a peak average cross-generator validation accuracy of $78.50\%$.
@@
 5.2 Fake Image Detection
-Following the experimental setup described in Section 4.2, we evaluate two methods: one based on neural networks (i.e., MLP) and the other on a $k$-NN approach.
+Following the experimental setup described in Section˜4.2, we evaluate two methods: one based on neural networks (i.e., MLP) and the other on a $k$-NN approach.
 Neural Network Classifier.
-We conduct preliminary experiments using the three MLPs described in Section 4.2, MLP-320, which achieves an average cross-generator accuracy of $79.17\%$, outperforms MLP-640 ($78.23\%$) and MLP640-320 ($78.94\%$).
-A detailed breakdown of the results for MLP-320 on cross-generator fake image detection is presented in Table 1.
-Trained on Tested on
+We conduct preliminary experiments using the three MLPs described in Section˜4.2.
+MLP-320, which achieves an average cross-generator accuracy of $79.1\%$, outperforms MLP-640 ($78.2\%$) and MLP-640-320 ($78.9\%$).
+A detailed breakdown of the results for MLP-320 on cross-generator fake image detection is presented in Table˜1.
+Trained on Accuracy (%) when tested on
 Midj. SDv1.4 SDv1.5 ADM Glide Wukong VQDM BigGAN Avg.
@@
 The results reveal a clear pattern: while the MLP effectively learns generator-specific artifacts (as evidenced by the high diagonal accuracy), it struggles to generalize to unseen generators.
-This behavior suggests the presence of distinct generator families that share underlying characteristics within the latent space of the diffusion U-Net.
-For instance, the MLP trained on Midjourney data performs well on images from SDV1.4, SDV1.5, and Glide, indicating a latent similarity among their generated outputs.
+We can also observe the presence of distinct generator families that share underlying characteristics within the latent space of the diffusion U-Net.
+For example, an MLP trained on SDM latents of Midjourney images transfers well to SDv1.4, SDv1.5, and Glide, suggesting a strong similarity in their latent feature representations.
 $k$-NN Classifier.
-While the features extracted from the diffusion U-Net are discriminative enough to separate real and synthetic data, the MLP classifier tends to learn specific artifacts of a given generator, which limits its ability to generalize.
-To address this limitation, we propose a $k$-NN approach.
-As a training-free method that relies on distances between feature prototypes, we hypothesize that $k$-NN can achieve a more robust generalization than the MLP.
-First, we identify the optimal $k$-NN configuration following the extensive parameter search described in Section 4.2.
-The model using the correlation distance, with $k=101$ and a support size of $2000$, achieved the best performance on the validation set (see Table 2).
-It yielded an average accuracy of $85.3\%$ across the eight generators, a significant improvement over the $79.17\%$ obtained by the best MLP approach.
-This optimized version of the $k$-NN model is finally tested on the GenImage test set (Table 3).
-Support set from Tested on
-Midj. SDV1.4 SDV1.5 ADM Glide Wukong VQDM BigGAN Avg.
-Midj. 88.1 87.0 86.9 84.7 88.3 88.3 87.3 86.3 87.1
-SDV1.4 88.3 89.9 89.6 83.2 86.7 62.6 83.8 88.3 84.0
-SDV1.5 89.4 89.7 90.4 84.4 89.8 87.1 87.2 88.7 88.3
-ADM 83.0 82.5 83.5 82.4 83.6 83.7 84.5 82.6 83.2
-Glide 85.7 83.9 84.1 85.4 87.4 87.4 87.8 83.8 85.7
-Wukong 81.2 75.5 76.7 84.5 89.5 88.6 87.8 77.8 82.7
-VQDM 83.0 83.4 84.1 83.6 84.5 84.7 85.5 83.0 84.0
-BigGAN 88.1 87.8 88.5 84.2 88.3 88.7 88.7 87.7 87.7
-Avg. 85.8 85.0 85.5 84.0 87.3 83.9 86.5 84.8 85.3
-Table 2: Cross-generator evaluation of $k$-NN on the GenImage validation set for fake image detection.
-We report the accuracy of the selected $k$-NN configuration.
-For each generator, we use the support set from one model and test it on the validation images of all the generators.
-Support set from Tested on
-Midj. SDV1.4 SDV1.5 ADM Glide Wukong VQDM BigGAN Avg.
-Midj. 91.4 89.1 89.3 88.0 91.5 91.6 89.9 88.6 89.9
-SDV1.4 89.9 90.9 90.7 85.4 88.2 64.1 84.7 89.4 85.4
-SDV1.5 91.7 91.7 91.4 86.7 91.9 89.9 89.4 90.9 90.4
-ADM 86.2 85.5 85.6 86.7 86.9 86.5 86.5 85.1 86.1
-Glide 89.0 86.7 86.8 89.6 91.1 90.6 90.0 87.2 88.9
-Wukong 85.3 78.7 79.5 88.2 93.0 92.7 91.2 81.6 86.3
-VQDM 86.7 86.6 86.5 87.9 88.6 88.1 87.9 85.8 87.3
-BigGAN 90.8 90.3 89.9 87.4 90.9 91.6 90.9 89.9 90.2
-Avg. 88.9 87.4 87.5 87.5 90.3 86.9 88.8 87.3 88.1
-Table 3: Cross-generator evaluation of $k$-NN on the GenImage test set for fake image detection.
-We report the accuracy of the selected $k$-NN configuration.
-For each generator, we use the support set from one model and test it on the images of all the generators.
+Although U-Net features effectively distinguish real from synthetic data, the MLP classifier tends to learn generator-specific artifacts, hindering generalization.
+To address this limitation, we propose a training-free $k$-NN approach that operates on distances between feature prototypes.
+By relying on the intrinsic geometry of the feature space rather than learned parameters, $k$-NN leverages the structure already encoded in the representations, leading to more effective cross-generator generalization.
+To this aim, guided by the hyperparameter search described in Section˜4.2, we select the $k$-NN model using correlation distance with $k=45$ and a size of $S_{G}=1000$.
+This configuration achieves the best validation accuracy, $85.3\%$ across eight generators as shown in Table˜2(a), and surpasses the best MLP ($79.1\%$).
+The same $k$-NN configuration is then evaluated on the GenImage test set (Table˜2(b)).
+(a) Results on GenImage Validation set
+$S_{G}$ Accuracy (%) when tested on Midj.
+SDv1.4 SDv1.5 ADM Glide Wukong VQDM BigGAN Avg.
+Midj.
+88.0 85.8 86.3 83.8 87.9 87.6 86.1 85.1 86.3 SDv1.4 88.1 90.0 90.7 83.5 87.7 65.9 84.7 88.1 84.8 SDv1.5 89.0 89.2 89.7 83.0 88.6 85.5 86.6 88.2 87.5 ADM 83.5 82.4 83.6 82.6 84.0 84.2 84.5 82.6 83.4 Glide 85.9 84.4 84.5 85.7 87.8 87.8 88.5 84.3 86.1 Wukong 82.1 74.8 75.7 84.5 89.7 89.0 88.5 76.8 82.6 VQDM 82.7 83.2 83.8 83.9 85.1 85.2 85.9 83.6 84.2 BigGAN 86.9 87.3 88.3 83.4 88.3 88.6 88.4 87.3 87.3 Avg.
+85.8 84.6 85.3 83.8 87.4 84.2 86.6 84.5 85.3
+(b) Results on GenImage Test set
+$S_{G}$ Accuracy (%) when tested on Midj.
+SDv1.4 SDv1.5 ADM Glide Wukong VQDM BigGAN Avg.
+Midj.
+91.1 88.4 88.7 87.3 91.0 90.8 88.8 88.6 89.3 SDv1.4 89.9 91.5 91.1 85.3 88.4 67.7 85.2 89.8 86.1 SDv1.5 91.6 91.2 91.1 85.8 91.0 87.9 88.8 90.3 89.7 ADM 86.4 85.7 85.9 87.0 87.1 86.8 86.9 85.2 86.4 Glide 89.3 87.0 86.9 89.8 91.4 90.8 90.4 87.5 89.1 Wukong 85.4 78.1 78.3 88.0 93.4 93.1 91.3 80.8 86.1 VQDM 86.9 86.7 86.7 88.3 89.0 88.6 88.5 86.1 87.6 BigGAN 89.8 89.6 89.5 86.3 90.5 91.5 90.3 89.5 89.6 Avg.
+88.8 87.3 87.3 87.2 90.2 87.2 88.8 87.2 88.0
+Table 2: Cross-generator evaluation of $k$-NN on the GenImage dataset for fake image detection. Accuracy of $k$-NN ($k=45$) classifier with support set size of 1,000 using distance correlation.
 Methods Tested on Avg.
 Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN
-CNNSpot [46] 58.2 70.3 70.2 57.0 57.1 67.7 56.7 56.6 61.7
-Spec [56] 56.7 72.4 72.3 57.9 65.4 70.3 61.7 64.3 65.1
-F3-Net [33] 55.1 73.1 73.1 66.5 57.8 72.3 62.1 56.5 64.6
-GramNet [26] 58.1 72.8 72.7 58.7 65.3 71.3 57.8 61.2 64.7
-DIRE [49] 65.0 73.7 73.7 61.9 69.1 74.3 63.4 56.7 67.2
-LaRE2 [29] 66.4 87.3 87.1 66.7 81.3 85.5 84.4 74.0 79.1
-LATTE [43] 71.3 79.3 81.8 82.2 92.8 82.0 82.9 87.8 82.5
-FRIDA (Ours) 88.9 87.5 87.5 87.5 90.3 86.9 88.9 87.4 88.1
-Table 4: Fake image detection comparison with state-of-the-art approaches on GenImage test set.
-The reported metric is the average percentage accuracy calculated from eight distinct models, each trained on the data from a different generator, and tested on the specified test set.
-† indicates that the results are reproduced by [29].
+CNNSpot [wang2020cnn] 58.2 70.3 70.2 57.0 57.1 67.7 56.7 56.6 61.7
+Spec [zhang2019detecting] 56.7 72.4 72.3 57.9 65.4 70.3 61.7 64.3 65.1
+F3-Net [qian2020thinking] 55.1 73.1 73.1 66.5 57.8 72.3 62.1 56.5 64.6
+GramNet [liu2020global] 58.1 72.8 72.7 58.7 65.3 71.3 57.8 61.2 64.7
+DIRE [wang2023dire] 65.0 73.7 73.7 61.9 69.1 74.3 63.4 56.7 67.2
+LaRE2 [luo2024lare] 66.4 87.3 87.1 66.7 81.3 85.5 84.4 74.0 79.1
+LATTE [vasilcoiu2025latte] 71.3 79.3 81.8 82.2 92.8 82.0 82.9 87.8 82.5
+FRIDA-200 (Ours) 86.9 86.2 86.1 85.0 87.9 87.5 86.9 85.9 86.5
+FRIDA (Ours) 88.8 87.3 87.3 87.2 90.2 87.2 88.8 87.2 88.0
+Table 3: Comparison with state-of-the-art methods for fake image detection on the GenImage test set.
+Given a specified test set, we evaluate eight $k$-NN classifiers—each based on the support set from one generator—and compute their average accuracy.
+FRIDA-200 exploits a support set of 200 samples, reduced by $80\%$ w.r.t.
+FRIDA.
 Best in bold, second best underlined.
-As detailed in Table 4, our proposed approach demonstrates a significant leap in performance over current state-of-the-art methods on the GenImage test set.
-The experimental setup is designed to rigorously test generalization: eight models are trained on limited data (2000 samples) from a single generator each, but evaluated on images from all generators.
+† results taken by [luo2024lare].
+As detailed in Table˜3, our approach demonstrates a significant leap in performance over current state-of-the-art methods on the GenImage test set.
+The experimental setup is designed to rigorously test generalization: eight models are trained on limited data (1,000 samples) from a single generator each, but evaluated on images from all generators.
 In this challenging scenario, our method is the one that better generalizes across these unseen data distributions, establishing a new state-of-the-art performance by a margin of nearly six percentage points.
-Beyond its accuracy, the primary advantage of our approach lies in its remarkable data and training efficiency.
-The strong results are achieved without a conventional training process; we use a pre–trained SDM and the $k$-NN framework merely requires storing the support set.
-The proposed approach maintains high performance even with a drastically smaller support set.
-When the sample size is reduced by $90\%$, the average cross-generator accuracy on the test set drops by only $1.6\%$ (from $88.1\%$ to $86.5\%$), a result that still outperforms the previous state-of-the-art (Table 4).
-This characteristic is crucial for practical deployment.
-In a field where new generative models are constantly emerging, our method provides a scalable and effective solution that does not require cost-prohibitive data generation and retraining cycles.
-Using an NVIDIA RTX 4090, SDM feature extraction required an average of 2.1 seconds, while $k$-NN classification took 0.0003 seconds.
-In contrast, on a CPU (Intel Core i7-9800X @ 3.80 GHz), feature extraction averaged 12.9 seconds and classification 0.002 seconds.
+Beyond its accuracy, the primary advantage of FRIDA in fake image detection lies in its remarkable data and training efficiency.
+The strong results are achieved without any conventional training process: we use a pre-trained SDM, and the $k$-NN framework merely requires storing the support set.
+The method also maintains high performance even with a drastically smaller support size.
+When the support set is reduced by $80\%$ (200 samples, referred to as FRIDA-200 in Table˜3), the average cross-generator accuracy on the test set drops by only $1.5\%$ (from $88.0\%$ to $86.5\%$), still outperforming the previous state-of-the-art (Table˜3).
+In a landscape where new generative models continually emerge, this property is crucial for practical deployment, as it avoids data-intensive retraining cycles.
 5.3 Source Model Attribution
-We compare MLP and $k$-NN classifiers on their ability to identify an image’s source generator, following the experimental procedure described in Section 4.3.
+We compare MLP and $k$-NN classifiers on their ability to identify an image’s source generator, following the experimental procedure described in Section˜4.3.
 $k$-NN Classifier.
-A portion of the results from the hyperparameter selection on the validation set obtained using the correlation distance is presented in Table 5.
-$k$ Support sizes 18 90 180 900 1800 3600 9000 1 25.2% 33.5% 36.4% 44.6% 48.1% 51.6% 55.8% 9 28.7% 31.3% 44.09% 48.7% 52.7% 57.7% 19 27.2% 42.5% 48.3% 52.1% 57.1% 35 39.5% 46.5% 50.9% 56.1% 101 39.7% 46.3% 52.0%
-Table 5: Effect of $k$ and support set size on source attribution accuracy.
-Validation accuracies for the nine-class attribution task (eight generators plus real images) computed across different values of $k$ and support set sizes.
-These results proved that the $k$-NN classifier is inadequate for effective source image attribution.
-Despite optimization (using a support set of 9,000 and k=9), its peak accuracy reached a mere $57.7\%$, a performance level far too low for practical application.
+We adapt the distance-based approach used for fake image detection for the task of source model attribution.
+A portion of the results from the hyperparameters selection on the validation set obtained using the correlation distance is presented in Table˜4.These results prove that the $k$-NN classifier is inadequate for effective source model attribution.
+Despite optimization (using a support set of 9,000 and $k=9$), its peak accuracy reached only $57.7\%$, a performance level far too low for practical application.
 Neural Network Classifier.
-We train the three MLP classifiers as described in Section 4.3; in Table 6 we report the average results of the MLPs on the validation set across ten training runs.
+We train the three MLP classifiers as described in Section˜4.3; in Table˜5 we report the average results of the MLPs on the validation set across 10 training runs.
+$k$ Support sizes 18 90 180 900 1800 3600 9000 1 25.2% 33.5% 36.4% 44.6% 48.1% 51.6% 55.8% 9 28.7% 31.3% 44.09% 48.7% 52.7% 57.7% 19 27.2% 42.5% 48.3% 52.1% 57.1% 35 39.5% 46.5% 50.9% 56.1% 45 37.8% 45.1% 50.2% 55.4%
+Table 4: Effect of $k$ and support set size on source model attribution accuracy. Validation accuracies for the nine-class attribution task.
 Model Avg. Accuracy Std.
@@
 MLP-640-320 84.78% 0.127
-Table 6: Source attribution performance of MLP models.
-Average accuracy and standard deviation over ten runs on the nine-class attribution task (eight generators plus real images) on the validation set.
-While the average accuracies of the three models on the source attribution task are comparable, the MLP-640, with $84.87\%$ of average accuracy, slightly outperforms the others and allows for an increase in performance of about 27 percentage points over the $k$-NN classifier.
+Table 5: Source model attribution performance of MLPs. Average validation accuracy and standard deviation over 10 runs.
+While the performance of the three models on the source model attribution task are comparable, the MLP-640, with $84.87\%$ of average accuracy, slightly outperforms the others and allows for an increase in performance of about 27 percentage points over the $k$-NN classifier.
 The MLP-640, when tested on the GenImage test set, achieves an accuracy of $84.36\%$.
-This result is consistent with the validation set performance, indicating a high degree of generalization.
-Figure 3: Confusion Matrix for source image attribution. The MLP-640 is evaluated on the GenImage test set.
-As shown in Figure 3, the classifier generally recognizes source generators with high accuracy; however, it struggles to distinguish models that share the same architecture.
-In particular, it frequently confuses SDM v1.4 with SDM v1.5, and, in some cases, the Wukong model — which is based on SDM — is also mistaken for them.
+This result is consistent with the validation performance, indicating a high degree of generalization.
+As shown in Figure˜3, the classifier generally recognizes source generators with high accuracy; however, it struggles to distinguish models that share the same architecture.
+Specifically, the model tends to confuse SDv1.4 with SDv1.5 and vice versa, while occasionally misattributing the SDM-based Wukong model to the same group.
+We attribute this phenomenon to their common architectural family.
+To validate this hypothesis, we conducted a family-based analysis by grouping data from these related sources, achieving a source model attribution accuracy of 96.67%.
 Notably, the huge performance increase of the MLP over the $k$-NN classifier on the validation set confirms that the features extracted by the diffusion U-Net contain generator-specific patterns.
-The MLP is capable of learning these distinguishing characteristics, whereas the $k$-NN approach, which relies solely on feature distances, cannot recognize these subtle patterns.
-Although $k$-NN proves effective for the broader task of distinguishing real from synthetic content, it is insufficient for the more nuanced task of source attribution.
-This suggests that while the feature distance between real and synthetic images is large, the distances between features from different generators are smaller.
-The key differentiators are not the absolute distances but rather specific, patterns within the features that the MLP can successfully identify.
+The MLP is capable of learning these distinctive characteristics, whereas the $k$-NN approach, which relies solely on feature distances, cannot recognize these subtle patterns.
+Although $k$-NN proves effective for the broader task of distinguishing real from synthetic content, it is insufficient for the more complex task of source model attribution.
+These results suggest that, while real and synthetic images can be detected through distances in the feature space, the discriminative power in source model attribution stems from more complex feature patterns that need an MLP to be captured.
 To investigate this hypothesis, we employ the SHAP algorithm to interpret the decisions of the MLP-640 classifier.
-We utilize the Gradient Explainer with a background dataset of 200 samples and 1000 test samples.
+We utilize the Gradient Explainer with a background dataset of 200 samples and 1,000 test samples.
 For each class, we identify the top 10 most informative features.
-In Figure 4 we report the percentage of the top 10 features shared between each pair of generators.
-Figure 4: Shared informative features across generators.
-We use SHAP with the Gradient Explainer to interpret the decisions of the MLP-640 classifier, employing 200 background and 1000 test samples.
-The plot reports, for each pair of generators, the percentage of overlap among their top 10 most informative features.
-[Real]Refer to caption [ADM]Refer to caption [BigGAN]Refer to caption
-[Glide]Refer to caption [Midjourney]Refer to caption [SDV1.4]Refer to caption
-[SDV1.5]Refer to caption [VQDM]Refer to caption [Wukong]Refer to caption
-Figure 5: Impact of the top-10 features on model decisions by SHAP analysis.
-Visualization of the ten most influential features identified by the SHAP analysis for each of the eight diffusion generators and the real image class.
-The figure highlights how different feature subsets contribute to the network’s decision process across generators, revealing both shared and model-specific attribution patterns.
+In Figure˜4 we report the percentage of the top 10 features shared between each pair of generators.
+Figure 3: Confusion Matrix for source model attribution. The MLP-640 is evaluated on the GenImage test set.
+Figure 4: Shared informative features across generators. Percentage of overlap among the top 10 SHAP features.
 The two SDMs exhibit a $50\%$ overlap in their top 10 most important features.
-Furthermore, these shared features influence the model’s output identically (see Figure 5 and Figure 5), indicating that the images generated by these two models lead to a latent representation with common characteristics.
+Furthermore, these shared features influence the model’s output identically (see LABEL:fig:shap_sdv14 and LABEL:fig:shap_sdv15 in the Supplementary Material), indicating that the images generated by these two models lead to a latent representation with common characteristics.
 As a result, the images generated by the two models are nearly indistinguishable to MLP.
 In contrast, the classifier successfully discriminates real and BigGAN images from the SDMs by exploiting divergent feature behavior.
-Although these models share $30\%$ of their top 10 features with the SDMs, the impact of these common features is markedly different (Figure 5 and Figure 5).
+Although these models share $30\%$ of their top 10 features with the SDMs, the impact of these common features is markedly different (LABEL:fig:shap_real and LABEL:fig:shap_BigGAN in the Supplementary Material).
 Finally, the Wukong model presents an interesting paradox.
-Despite occasional misclassifications with the SDM-generated images, the SHAP analysis reveals zero feature overlap within the top 10 most influential features (Figure 5).
+Despite occasional misclassifications with the SDM-generated images, the SHAP analysis reveals zero feature overlap within the top 10 most influential features (LABEL:fig:shap_wukong in the Supplementary Material).
 This indicates that the classifier’s errors are not driven by the same high-impact features that define the SDM class.
 Instead, the confusion likely arises because a different set of features in Wukong images produces a combined effect that coincidentally mimics the characteristics of an SDM, leading the classifier to an incorrect conclusion.
+5.4 Extended Experimental Analysis
+Backbone Comparison.
+In Table˜6 we report the results of the comparison among SDM, CLIP, and DINO as feature extractors following the setup described in Section˜4.4.
+Features from SDM consistently outperform features from the ViT encoders, validating the assumption that diffusion-based features are inherently more effective at isolating the discriminative cues necessary for distinguishing real from generated imagery.
+Robustness to Image Perturbations.
+In Figure˜5 we evaluate robustness under JPEG compression, Gaussian noise and Gaussian blur.
+While accuracy remains stable for JPEG factors up to 30 (a), it shows a steady decay under Gaussian noise (b).
+Notably, Gaussian blur (c) induces a sharp initial drop at $\sigma=1$ followed by a plateau, suggesting that the model successfully pivots to structural cues when local textures are suppressed.
+Out-of-Distribution Robustness.
+To evaluate the generalization capability of our method, we test FRIDA on several OOD datasets.
+Our approach achieves a solid average accuracy of 73.5%, demonstrating consistent detection across diverse generative architectures.
+Specifically, the model yields 69.4% on Flux 1.0 Dev, 68.1% on Flux 1.1 Pro, 72.5% on GPT Image 1, and 73.0% on SDv3.5.
+We observe a significant performance peak on Qwen (95.3%), while even on the most challenging generator in our benchmark, Ideogram 3.0, the detector maintains an accuracy of 62.7%.
+Figure 5: Robustness analysis against image distortions. We report the detection accuracy for different image perturbations: (a) JPEG compression, (b) Gaussian noise, and (c) Gaussian blur.
+Backbone Midj. SDv1.4 SDv1.5 ADM GLIDE BigGAN VQDM Wukong Avg
+SD (Ours) 88.8 87.3 87.3 87.2 90.2 87.2 88.8 87.2 88.0
+CLIP 63.6 64.6 64.5 57.9 66.3 60.2 55.9 64.1 61.5
+DINO 58.0 62.5 62.3 57.5 66.2 63.0 61.9 63.7 61.2
+Table 6: Backbone comparison. $k$-NN fake image detection replacing SDM with ViT-based backbones (CLIP and DINO).
 6 Conclusions
-In this work, we presented FRIDA, a training-free framework that repurposes internal activations of pre-trained diffusion models for deepfake detection and source generator attribution.
-FRIDA operates entirely in the diffusion feature space, leveraging the latent representation of a Stable Diffusion U-Net as discriminative descriptors of image authenticity.
-A simple k-nearest-neighbor classifier applied to these features achieves state-of-the-art detection performance and generalizes effectively across unseen generators, while a compact neural model trained on the same representations enables accurate source attribution.
-Our analysis shows that diffusion features, although optimized for image generation, are surprisingly effective for fake image detection and source model attribution.
-This finding suggests that diffusion features can serve as a universal and interpretable basis for synthetic image forensics, bridging generative modeling and authenticity analysis.
+In this work, we introduced FRIDA, a framework that leverages the internal representation of a pre-trained diffusion model for fake image detection and source model attribution.
+For fake image detection, FRIDA entirely operates within the diffusion feature space, exploiting latent representations from the Stable Diffusion U-Net as compact and discriminative descriptors of image authenticity.
+Through extensive evaluation, we demonstrated that a simple $k$-nearest neighbour classifier applied to these features achieves state-of-the-art detection performance with a limited support set, generalizing effectively to unseen generators without any retraining.
+For source model attribution, a compact neural classifier trained on the same features accurately identifies the generative source, confirming that diffusion features encode model-specific fingerprints.
+Furthermore, we show that diffusion-based features are more discriminative than CLIP and DINO features for fake image detection.
+Through extensive testing, we also demonstrate that our method maintains high performance under various image distortions and effectively generalizes to out-of-distribution generators, including state-of-the-art models like Flux and Qwen.
+These findings may position diffusion representations as a reliable foundation for AI-generated image forensics.